OpenStudy (anonymous):

write a recursive definition for the sequence 8,6,4,2,...?

OpenStudy (anonymous):

It is an A.P first term a=8 common difference d=any term-previous term=6-8=-2 n th term tn=a+(n-1)d plug the values of a and d and find tn

OpenStudy (anonymous):

so it would be a1=8; an=a^n-1-2?

OpenStudy (anonymous):

i thik it would be\[a _{n}=10-2n\] with n=1,2,3,4..........

OpenStudy (anonymous):

tn or an=8+(n-1)*-2=8-2n+2=10-2n